# Motives of nullcones of quiver representations

### v1

@@ 4.1. Introduction
(2) $\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in L_{\mathbf{d}}}\mathbb{L}^{\phi(s)}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}~{}.$
@@
with stability $\theta_{s}=(-1,1)$ and $\theta_{s}=(0)$, respectively.
@@ 4.2. Rational/integral level quiver
We know that the elements $s$ of the index set $\tilde{L}_{\mathbf{d}}$ gives rise to quivers $Q_{s}$, such that the corresponding Hesselink stratum is non-empty if and only if the semi-stable locus is empty. Our next aim is now to define a so-called rational level quiver which combines the level quivers associated to the segments, in such a way that the previous properties are still preserved, but the index set for the Hesselink stratification can be described in terms of specific dimension vectors $\mathbf{e}$ for this fixed quiver.
@@
for every arrow $\alpha:i\rightarrow j$ in $Q$ and every $a\in\mathbb{Q}$. This quiver obviously admits a decomposition into components $Q_{\mu}$ for $\mu\in[0,1)\cap\mathbb{Q}$, where $Q_{\mu}$ is the full subquiver of all vertices $(i,a)$ such that the fractional part $\{a\}$ of $a$ equals $\mu$.
We define a stability on $Q_{\mathbb{Q}}$ by $\theta(i,a)=-a$. We call a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ balanced if $\theta(\mathbf{e}|_{Q_{\mu}})=0$ for all $\mu\in[0,1[\cap\mathbb{Q}$ (where $\mathbf{e}|_{Q_{\mu}}$ denotes the restriction of $\mathbf{e}$ to the component $Q_{\mu}$). Every finite dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ admits a projection $|\mathbf{e}|$ to a dimension vector for $Q$ by $|\mathbf{e}|_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}.$
@@
Using this notation, the index set $\tilde{L}_{\mathbf{d}}$ for the Hesselink stratification of $N_{\mathbf{d}}(Q)$ is the set of all balanced dimension vectors $\mathbf{e}$ for $Q_{\mathbb{Q}}$ such that $|\mathbf{e}|=\mathbf{d}$. Moreover, we have an isomorphism of semistable loci
@@
For every entry $e_{i,a}$ in $\mathbf{e}$ we take $e_{i,a}$-times the entry $a$ in the coweight, in weakly ascending order for every vertex of $Q$. That this induces a bijection between balanced dimension vectors and balanced coweights now follows immediately from the definitions.
@@
We denote the set of balanced dimension vectors $\mathbf{e}$ for $\mathbb{Q}_{\mathbb{Q}}$ such that $|{\bf e}|={\bf d}$ by $L^{\prime}_{d}$.
The quiver $Q_{\mathbb{Q}}$ is the disjoint union of the $Q_{\mu}$ for $\mu\in[0,1[\cap\mathbb{Q}$, and every $Q_{\mu}$ can be identified with $Q_{0}=Q_{\mathbb{Z}}$, called the integral level quiver. We can reparametrize the index set $L^{\prime}_{\bf d}$: we map $\mathbf{e}$ to the tuple $(\mathbf{e}|_{Q_{\mu}})_{\mu}$ of dimension vectors, which we view as dimension vectors for $Q_{\mathbf{Z}}$. Now we interprete this reparametrization in terms of our stability conditions:
@@
$\sigma(\mathbf{e}_{\mu})=\frac{\theta(\mathbf{e}_{\mu})}{\dim\mathbf{e}_{\mu}}=-\mu~{}$
@@
The index set $L^{\prime}_{\bf d}$ for the Hesselink stratification can be viewed as the set of tuples $(\mathbf{e}_{\mu})_{\mu}$ of dimension vectors for $Q_{\mathbf{Z}}$ such that ${\bf e}_{\mu}=0$ or $\sigma(\mathbf{e}_{\mu})=-\mu$ for all $\mu$, and such that $\sum_{\mu}|\mathbf{e}_{\mu}|=\mathbf{d}.$ We will denote the set of such tuples by $L$ in the following.
@@
The index sets $L^{\prime}_{\bf d}$ and $L$ are obviously in bijection by the previous construction. Similarly to the proof of Proposition4.9, the semistable locus does not change when passing from $\theta$ to $\theta-y\cdot\dim$ for an arbitrary $y$. ∎
@@ 5. Wall-crossing type formula
Using Theorem 4.6, we see that $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$ is the dimension of the space of $|Q_{1}|$-tuples of $d\times d$-matrices with zero entries except possibly at position $(l,k)$ in the matrix corresponding to $\alpha:i\rightarrow j^{\prime}$ if $k\in I_{i}$, $l\in I_{j}$ and $s_{l}-s_{k}>1$. In terms of the rational level quiver this dimension is encoded in the dimensions $e_{i,a}$ at vertices which are not directly adjacent:
@@
Since we have shifted the dimension vector $\mathbf{e}$ to a tuple $(\mathbf{e}_{\mu})_{\mu}$ of dimension vectors for $Q_{\mathbb{Z}}$, we can rewrite this formula in terms of $(\mathbf{e}_{\mu})_{\mu}$. In the first summand, we replace $e_{i,a}$ and $e_{j,b}$ by $e_{i,a-\mu}^{\mu}$ and $e_{j,b-\mu^{\prime}}^{\mu^{\prime}}$, respectively, and similarly in the second term:
@@
With our new parametrization of the index set and propositions 4.9 and 4.10, we can now state the following formula for the motive of $N_{\mathbf{d}}$:
@@
$=\sum_{(\mathbf{e}_{\mu})_{\mu}\in L}\mathbb{L}^{\phi((\mathbf{e}_{\mu})_{\mu})}\prod_{\mu}\frac{[R_{\mathbf{e}_{\mu}}^{sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}~{}.$
@@
We define the $k$-shift ${\sigma}^{k}\mathbf{e}$ of a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Z}}$ as the dimension vector
$({\sigma}^{k}\mathbf{e})_{i,a}=\mathbf{e}_{i,a+k}~{}.$
@@
$-\langle\mathbf{e}^{\mu},{\sigma}^{k}\mathbf{e}^{\mu^{\prime}}\rangle=-\sum_{i}\sum_{a}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{i\rightarrow j}\sum_{a}e_{i,a}^{\mu}e_{j,a+1+k}^{\mu^{\prime}}~{}.$
and we can now rewrite $\phi((\mathbf{e}_{\mu})_{\mu})$, which equals
@@
$-\sum_{i}(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}),$
as
@@
$-\sum_{i}\sum_{a\in\mathbb{Z}}(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,a}^{\mu^{\prime}})$
@@
$-\sum_{k>0}\sum_{\mu,\mu^{\prime}}\langle\mathbf{e}_{\mu},{\sigma}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle~{}.$
@@
$\sum\mathbb{L}^{-\sum_{k>0}\sum_{\mu,\mu^{\prime}\in[0,1[\cap\mathbb{Q}}\langle\mathbf{e}_{\mu},{\sigma}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle}\prod_{\mu\in[0,1[\cap\mathbb{Q}}\frac{[R_{\mathbf{e}_{\mu}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}t^{|\sum_{\mu}\mathbf{e}_{\mu}|}~{},$
where the sum ranges over all tuples $({\bf e}_{\mu})_{\mu\in[0,1[\cap\mathbb{Q}}$ of dimension vectors for $Q_{\mathbb{Z}}$ such that ${\bf e}_{\mu}=0$ for all but finitely many $\mu$, and $\sigma({\bf e}_{\mu})=-\mu$ otherwise.
@@
$u^{\mathbf{e}}\ast u^{\mathbf{e}^{\prime}}\coloneq\mathbb{L}^{-\sum_{k\geq 0}\langle\mathbf{e},{\sigma}^{k}\mathbf{e}^{\prime}\rangle_{\mathbb{Q}_{\mathbb{Z}}}-\sum_{k>0}\langle\mathbf{e}^{\prime},{\sigma}^{k}\mathbf{e}\rangle_{\mathbb{Q}_{\mathbb{Z}}}}\cdot u^{\mathbf{e}+\mathbf{e}^{\prime}}~{}.$
@@
$\prod^{\rightarrow}_{\mu\in[0,1[\cap\mathbb{Q}}\Biggl{(}\sum\limits_{\begin{subarray}{c}\mathbf{e}\in\mathbb{N}(Q_{\mathbb{Z}})_{0}\\ \operatorname{supp}(\mathbf{e})\text{ finite}\end{subarray}}\mathbb{L}^{-\sum_{k>0}\langle\mathbf{e},{\sigma}^{k}\mathbf{e}\rangle}\frac{[R_{\mathbf{e}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}}]}\cdot u^{\mathbf{e}}\Biggr{)}~{}.$

### v2

@@ 4.1. Introduction
(2) $\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\tilde{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}~{}.$
Now we extend the index set $\tilde{L}_{\mathbf{d}}$ of dominant and minimal string to the index set $\bar{L}_{\mathbf{d}}$ of dominant and balanced strings. The following theorem shows that this change of index sets does not affect the formula of the motive of the nullcone.
Theorem 4.7.
For a balanced string $\mu$ which is not minimal, the semistable locus $R_{d_{s}}^{\theta_{s}-sst}(Q_{s})$ is empty.
Proof.
For a segment $K=J_{i}$, we define $Q_{K}=Q_{i}$ and ${\bf d}_{K}={\bf d}_{i}$. First note that we can switch from the stability $\theta_{s}$ to the stability $\tilde{\theta}$ given by $\tilde{\theta}(v,j)=-(p_{i}+j)$ on the level quiver $Q_{K}$ for a segment $K$ starting with $p_{i}$, since the open subset of semistable representations does not change when passing from $\theta$ to $x\cdot\theta$ for a positive $x$. We also have $\tilde{\theta}({\bf d}_{K})=0$ for all segments $K$
Moreover, we note that for a fixed string $s=(s_{1},\dots,s_{n})\in\mathbb{Q}^{n}$, the minimum of $\{|s-q|:=\sum_{i=1}^{n}(s_{i}-q)^{2}\}$ for $q\in\mathbb{Q}$ is given by $|s-\bar{q}|$ such that $\sum_{i=1}^{n}(s_{i}-\bar{q})=0$. If, additionally, $\sum_{i=1}^{n}s_{i}<\sum_{i=1}^{n}(s_{i}-q)<0$, then $|s|\geq|s-q|$, since we can consider $|s-q|$ (for a fixed $s$) as a quadratic function in $q$. Analogously, if $\sum_{i=1}^{n}s_{i}>\sum_{i=1}^{n}(s_{i}-q)>0$, then $|s|\geq|s-q|$.
Let $\mu=(\mu_{1},\dots,\mu_{d})$ be a balanced, but not minimal string. By definition, there exists a (balanced) $\mu^{\prime}$ with $|\mu^{\prime}|<|\mu|$ and $R(\mu^{\prime})\supset R(\mu)$. Therefore we find a segment $J=(\mu_{i})_{i\in I}$ in $\mu$ such that the corresponding substring $J^{\prime}=(\mu^{\prime})_{i\in I}$ of $\mu^{\prime}$ satisfies $|J^{\prime}|=\sum_{I}\mu_{i}^{\prime 2}<\sum_{I}\mu_{i}^{2}=|J|$. We can decompose the indexing set $I$ in three disjoint sets given by
$I^{n}_{1}=\{i\in I\text{ with }\mu_{i}>\mu_{i}^{\prime}\}~{},$
$I^{p}_{1}=\{i\in I\text{ with }\mu_{i}<\mu_{i}^{\prime}\}~{},$
$I_{0}=\{i\in I\text{ with }\mu_{i}=\mu_{i}^{\prime}\}~{}.$
Since $|J^{\prime}|<|J|$, either $\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ or $\sum_{I^{p}_{1}}\mu_{i}^{\prime 2}<\sum_{I^{p}_{1}}\mu_{i}^{2}$. Without loss of generality we assume the inequality holds for $I^{n}_{1}$. The other case follows analogously.
We call a subset $\hat{I}$ of $I$ left-complete if, for every $k\in\hat{I}$, all $l$ with $\mu_{k}-\mu_{l}=1$ and $\pi_{lk}\in\Pi$ also belong to $\hat{I}$. Note that this implies the existence of an arrow from the vertex corresponding to $\mu_{l}$ to the vertex corresponding to $\mu_{k}$ in $Q_{J}$.
Consider $\mu_{i}^{\prime}$ with $i\in I^{n}_{1}$. Then there exists an $\epsilon_{i}>0$ with $\mu_{i}^{\prime}=\mu_{i}-\epsilon_{i}$. In the level quiver $Q_{J}$, there exists a vertex $v_{i}$ corresponding to $\mu_{i}$. Every arrow $\alpha$ in $Q_{J}$ with $t(\alpha)=v_{i}$ corresponds to (possibly several) elements $\mu_{l}\in J$ such that $\mu_{i}-\mu_{l}=1$ and $\pi_{li}\in\Pi$. Since $R(\mu^{\prime})\supset R(\mu)$, every $\mu_{l}$ obtained in the above way fulfills $\mu^{\prime}_{l}=\mu_{l}-\epsilon_{l}$ with $\epsilon_{l}>\epsilon_{i}$. This shows that $I^{n}_{1}$ is left-complete.
For a subset $\hat{I}$ of $I$ we can now construct a subquiver $\hat{Q}_{J}$ of $Q_{J}$ from the substring $\hat{J}$ of $J$ determined by $\hat{I}$ in the same way the level quiver is constructed from a balanced string; we can also construct a dimension vector $\hat{{\bf d}}_{J}$ for this subquiver which fulfills $(\hat{{\bf d}}_{J})_{i}\leq({\bf d}_{J})_{i}$.
We now assume that the semistable locus in question is not empty. We then claim that for all left-complete subsets $\hat{I}$, the inequality $\sum_{\hat{I}}\mu_{i}\leq 0$ holds. Assume, to the contrary, that $\sum_{\hat{I}}\mu_{i}>0$. Let $V$ be an arbitrary representation of dimension vector ${\bf d}_{J}$ of $Q_{J}$. Since $\hat{I}$ is left-complete,we have $({\bf d}_{J}-\hat{\bf d}_{J})_{k}=0$ for all vertices $v_{k}$ in $\hat{Q}_{J}$, which are not a sink in $\hat{Q}_{J}$, and $({\bf d}_{J}-\hat{\bf d}_{J})_{k}=({\bf d}_{J})_{k}$ for all $v_{k}$ which are not in $\hat{Q}_{J}$. Therefore, $V$ admits a subrepresentation $V^{\prime}$ of dimension vector ${\bf d}_{J}-\hat{\bf d}_{J}$. We have $\tilde{\theta}(d_{J})=0$ since $\mu$ is balanced, showing that $\tilde{\theta}(d_{J}-\hat{d}_{J})>0$. This proves the claim.
We know that $\sum_{I^{n}_{1}}\mu_{i}^{\prime 2}=\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$ and $\sum_{I^{n}_{1}}\mu_{i}\leq 0$, because $I^{n}_{1}$ is left-complete. We choose $\tilde{\epsilon}_{1}$ as the minimum of all $\epsilon_{i}$ for $i\in I^{n}_{1}$. We get
$\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}<\sum_{I^{n}_{1}}\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}\leq\sum_{I^{n}_{1}}\mu_{i}\leq 0~{},$
leading to $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}$. Now we define
$I^{n}_{2}=\{i\in I^{n}_{1}:\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}<\mu_{i}\}~{}.$
In other words, we exclude all $\mu_{i}^{\prime}$ with $\mu_{i}^{\prime}+\tilde{\epsilon}_{1}=\mu_{i}$. Now $I^{n}_{2}$ is left-complete by the same argument as for $I^{n}_{1}$. We choose $\tilde{\epsilon}_{2}$ as the minimum of all $\epsilon_{i}-\tilde{\epsilon}_{1}$ with $i\in I^{n}_{2}$. We get:
$\sum_{I^{n}_{2}}\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}<\sum_{I^{n}_{2}}\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2}\leq\sum_{I^{n}_{2}}\mu_{i}\leq 0$
concluding
$\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}~{}.$
We get a descending chain of subsets
$I_{1}^{n}\supsetneq I_{2}^{n}\supsetneq\dots~{},$
which has to terminate after finitely many steps. This yields the following estimate: $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}>$
$>\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1})^{2}>\sum_{I^{n}_{1}\setminus I^{n}_{2}}\mu_{i}^{2}+\sum_{I^{n}_{2}}(\mu_{i}-\epsilon_{i}+\tilde{\epsilon}_{1}+\tilde{\epsilon}_{2})^{2}>\ldots>\sum_{I^{n}_{1}}\mu_{i}^{2}~{}.$
Since we have started with $\sum_{I^{n}_{1}}(\mu_{i}-\epsilon_{i})^{2}<\sum_{I^{n}_{1}}\mu_{i}^{2}$, we have thus constructed a contradiction, proving that a balanced, but not minimal string, has an empty semistable locus.
Corollary 4.8.
For the nullcone $N_{\bf d}$ we have the following equation of motives:
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in{\bar{L}_{\mathbf{d}}}}\mathbb{L}^{\phi(s)}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}~{}.$
@@
with stability $\theta_{s}=(1,-1)$ and $\theta_{s}=(0)$, respectively.
@@ 4.2. Rational/integral level quiver
We know that the elements $s$ of the index set ${\bar{L}_{\mathbf{d}}}$ give rise to quivers $Q_{s}$, such that the corresponding Hesselink stratum is non-empty if and only if the semi-stable locus is non-empty. Our next aim is now to define a so-called rational level quiver which combines the level quivers associated to the segments, in such a way that the previous properties are still preserved, but the index set for the Hesselink stratification can be described in terms of specific dimension vectors $\mathbf{e}$ for this fixed quiver.
@@
for every arrow $\alpha:i\rightarrow j$ in $Q$ and every $a\in\mathbb{Q}$. This quiver obviously admits a decomposition into components $Q_{\mu}$ for $\mu\in[0,1[\cap\mathbb{Q}$, where $Q_{\mu}$ is the full subquiver of all vertices $(i,a)$ such that the fractional part $\{a\}$ of $a$ equals $\mu$.
We define a stability on $Q_{\mathbb{Q}}$ by $\theta(i,a)=-a$. We call a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ balanced if $\theta(\mathbf{e}|_{Q_{\mu}})=0$ for all $\mu\in[0,1[\cap\mathbb{Q}$ (where $\mathbf{e}|_{Q_{\mu}}$ denotes the restriction of $\mathbf{e}$ to the component $Q_{\mu}$). Every finitely supported dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ admits a projection $|\mathbf{e}|$ to a dimension vector for $Q$ by $|\mathbf{e}|_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}.$
@@
Using this notation, the index set ${\bar{L}_{\mathbf{d}}}$ for the Hesselink stratification of $N_{\mathbf{d}}(Q)$ is in bijection to the set of all balanced dimension vectors $\mathbf{e}$ for $Q_{\mathbb{Q}}$ such that $|\mathbf{e}|=\mathbf{d}$. Moreover, we have an isomorphism of semistable loci
@@
For every entry $e_{i,a}$ in $\mathbf{e}$ we take $e_{i,a}$-times the entry $a$ in the coweight, in weakly ascending order for every vertex of $Q$. That this induces a bijection between balanced dimension vectors and balanced and dominant coweights now follows immediately from the definitions.
@@
We denote the set of balanced dimension vectors $\mathbf{e}$ for ${Q}_{\mathbb{Q}}$ such that $|{\bf e}|={\bf d}$ by $L^{\prime}_{\mathbf{d}}$.
The quiver $Q_{\mathbb{Q}}$ is the disjoint union of the $Q_{\mu}$ for $\mu\in[0,1[\cap\mathbb{Q}$, and every $Q_{\mu}$ can be identified with $Q_{\mu=0}=Q_{\mathbb{Z}}$, called the integral level quiver. We can reparametrize the index set $L^{\prime}_{\bf d}$: we map $\mathbf{e}$ to the tuple $(\mathbf{e}|_{Q_{\mu}})_{\mu}$ of dimension vectors, which we view as dimension vectors for $Q_{\mathbb{Z}}$. Now we interprete this reparametrization in terms of our stability conditions:
@@
$\sigma(\mathbf{e}_{\mu})=\frac{\theta(\mathbf{e}_{\mu})}{\dim\mathbf{e}_{\mu}}=\mu~{}$
@@
The index set $L^{\prime}_{\bf d}$ for the Hesselink stratification can be viewed as the set of tuples $(\mathbf{e}_{\mu})_{\mu}$ of dimension vectors for $Q_{\mathbb{Z}}$ such that ${\bf e}_{\mu}=0$ or $\sigma(\mathbf{e}_{\mu})=\mu$ for all $\mu$, and such that $\sum_{\mu}|\mathbf{e}_{\mu}|=\mathbf{d}.$ We will denote the set of such tuples by $L({\bf d})$ in the following.
@@
The index sets $L^{\prime}_{\bf d}$ and $L({\bf d})$ are obviously in bijection by the previous construction. Similarly to the proof of Proposition4.11, the semistable locus does not change when passing from $\theta$ to $\theta-y\cdot\dim$ for an arbitrary $y$. ∎
@@ 5. Wall-crossing type formula
Using Theorem 4.6, we see that $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$ is the dimension of the space of $|Q_{1}|$-tuples of $d\times d$-matrices with zero entries except possibly at position $(l,k)$ in the matrix corresponding to $\alpha:i\rightarrow j$ if $k\in I_{i}$, $l\in I_{j}$ and $s_{l}-s_{k}>1$. In terms of the rational level quiver this dimension is encoded in the dimensions $e_{i,a}$ at vertices which are not directly adjacent:
@@
Since we have shifted the dimension vector $\mathbf{e}$ to a tuple $(\mathbf{e}_{\mu})_{\mu}$ of dimension vectors for $Q_{\mathbb{Z}}$, we can rewrite this formula in terms of $(\mathbf{e}_{\mu})_{\mu}$. In the first summand, we replace $e_{i,a}$ and $e_{j,b}$ by $e_{i,a-\mu}^{\mu}$ and $e_{j,b-\mu^{\prime}}^{\mu^{\prime}}$, respectively, and similarly in the second term. Note that, consequently, $a$ and $b$ denote integers in all the following summations:
@@
With our new parametrization of the index set and propositions 4.11 and 4.12, we can now state the following formula for the motive of $N_{\mathbf{d}}$:
@@
$=\sum_{(\mathbf{e}_{\mu})_{\mu}\in L}\mathbb{L}^{\phi((\mathbf{e}_{\mu})_{\mu})}\prod_{\mu}\frac{[R_{\mathbf{e}_{\mu}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}~{}.$
@@
We define the $k$-shift ${\tau}^{k}\mathbf{e}$ of a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Z}}$ as the dimension vector
$({\tau}^{k}\mathbf{e})_{i,a}=\mathbf{e}_{i,a+k}~{}.$
@@
$-\langle\mathbf{e}^{\mu},{\tau}^{k}\mathbf{e}^{\mu^{\prime}}\rangle=-\sum_{i}\sum_{a}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{i\rightarrow j}\sum_{a}e_{i,a}^{\mu}e_{j,a+1+k}^{\mu^{\prime}}~{}.$
We now consider $\phi((\mathbf{e}_{\mu})_{\mu})$, which equals
@@
$-\sum_{i}(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}).$
We rewrite this as
@@
$-\sum_{i}\sum_{a\in\mathbb{Z}}(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,a}^{\mu^{\prime}}),$
@@
$-\sum_{k>0}\sum_{\mu,\mu^{\prime}}\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle~{}.$
@@
$\sum\mathbb{L}^{-\sum_{k>0}\sum_{\mu,\mu^{\prime}\in[0,1[\cap\mathbb{Q}}\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle}\prod_{\mu\in[0,1[\cap\mathbb{Q}}\frac{[R_{\mathbf{e}_{\mu}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}t^{|\sum_{\mu}\mathbf{e}_{\mu}|}~{},$
where the sum ranges over all sets $L({\bf d})$, that is, over all tuples $({\bf e}_{\mu})_{\mu\in[0,1[\cap\mathbb{Q}}$ of dimension vectors for $Q_{\mathbb{Z}}$ such that ${\bf e}_{\mu}=0$ for all but finitely many $\mu$, and $\sigma({\bf e}_{\mu})=\mu$ otherwise.
@@
$u^{\mathbf{e}}\ast u^{\mathbf{e}^{\prime}}=\mathbb{L}^{-\sum_{k\geq 0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}^{\prime}\rangle_{\mathbb{Q}_{\mathbb{Z}}}-\sum_{k>0}\langle\mathbf{e}^{\prime},{\tau}^{k}\mathbf{e}\rangle_{\mathbb{Q}_{\mathbb{Z}}}}\cdot u^{\mathbf{e}+\mathbf{e}^{\prime}}~{}.$
@@
$\prod^{\rightarrow}_{\mu\in[0,1[\cap\mathbb{Q}}\Biggl{(}\sum\limits_{\begin{subarray}{c}\mathbf{e}\in\mathbb{N}(Q_{\mathbb{Z}})_{0}\\ \operatorname{supp}(\mathbf{e})\text{ finite}\end{subarray}}\mathbb{L}^{-\sum_{k>0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}\rangle}\frac{[R_{\mathbf{e}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}}]}\cdot u^{\mathbf{e}}\Biggr{)}~{}.$